# Learning with pyCub: A New Simulation and Exercise Framework for Humanoid Robotics

### v1

@@ III-E Grasping
IV Conclusion, Discussion, and Future Work
We presented pyCub, an unified educational platform consisting of a physics-based simulation of the humanoid robot iCub and exercises that aim to to teach core robotic tasks and skills. The whole framework is open-source and available at https://rustlluk.github.io/pyCub. The exercises cover basic robot control in joint and Cartesian space, more complicated Cartesian movements, gazing at a moving target, reactive movements in velocity space to avoid obstacles detected using signal processing from whole-body artificial skin, and grasping based on computer vision techniques.
The simulation and exercises were deployed in two runs of a course on humanoid robotics. Previous runs were taught using iCub’s original Gazebo simulator controlled through a middleware YARP [21] using C++. Switching to pyCub, a Python-based simulator changed the dynamics of the class. Although the students were engineering students with experience with C++, the whole system was troublesome to learn and most complaints in the post-semester survey were about YARP and C++. Using pyCub allowed them to focus more on the robotic aspects of the exercises rather than on new middleware or compilation issues. For example, although the first exercise is mainly intended to get used to the simulator, the students showed that it is also an engaging exercise for them. The goal is to get the ball as far as possible from a table, and the expected solution is to simply hit the ball with an arm. However, there were many solutions such as grasping and throwing or even kicking the ball—see the accompanying video or the website of the project at https://rustlluk.github.io/pyCub.

### v2

@@ 3.5 Grasping
4 Conclusion, Discussion, and Future Work
We presented pyCub, an unified educational platform consisting of a physics-based simulation of the humanoid robot iCub and exercises that aim to teach core robotic tasks and skills. The whole framework is open-source and available at https://rustlluk.github.io/pyCub. The exercises cover basic robot control in joint and Cartesian space, more complicated Cartesian movements, gazing at a moving target, reactive movements in velocity space to avoid obstacles detected using signal processing from whole-body artificial skin, and grasping based on computer vision techniques.
The simulation and exercises were deployed in two runs of a course on humanoid robotics. Previous runs were taught using iCub’s original Gazebo simulator controlled through a middleware YARP [15] using C++. Switching to pyCub, a Python-based simulator changed the dynamics of the class. Although the students were engineering students with experience with C++, the whole system was troublesome to learn and most of the complaints in the post-semester survey were about YARP and C++. Using pyCub allowed them to focus more on the robotic aspects of the exercises rather than on new middleware or compilation issues. For example, although the first exercise is mainly intended to get used to the simulator, the students showed that it is also an engaging exercise for them. The goal is to get the ball as far as possible from a table, and the expected solution is to simply hit the ball with an arm. However, there were many solutions such as grasping and throwing or even kicking the ball—see the accompanying video or the website of the project at https://rustlluk.github.io/pyCub.
@@ 4 Conclusion, Discussion, and Future Work
Acknowledgments
This work was co-funded by the European Union under the project Robotics and Advanced Industrial Production (reg. no. CZ.02.01.01/00/22_008/0004590). L.R. was additionally supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS24/096/OHK3/2T/13.